Take the Laplace transform of the following initial value problem and solve for Y(s)=L{y(t)}:

y′′−5y′−24y=S(t) y(0)=0,y′(0)=0   

Where S(t)={1, ,0≤t<1 0, 1≤t<2}    S(t+2)=S(t)

Y(s) = ?
Answer #1

y-541–244 = SCE) LESCE)) = ? vos zo L{SCE)} = ? -400 =0 SCL) = CUOCE)- L{506) = 2005 .. Lcy) - 52(41)–24 Lcy = L&sity] slycfirst we find laplace transform of s(t) using heavy side function. ua(t) is the heavy side function with step size a.

L{ua(t)}=e-as/s